# UniCalli: A Unified Diffusion Framework for Column-Level Generation and Recognition of Chinese Calligraphy

### v1

@@ Abstract
Computational replication of Chinese calligraphy remains challenging. Existing methods falter, either creating high-quality isolated characters while ignoring page-level aesthetics like ligatures and spacing, or attempting page synthesis at the expense of calligraphic correctness. We introduce UniCalli, a unified diffusion framework for column-level recognition and generation. Training both tasks jointly is deliberate: recognition constrains the generator to preserve character structure, while generation provides style and layout priors. This synergy fosters concept-level abstractions that improve both tasks, especially in limited-data regimes. We curated a dataset of over 8,000 digitized pieces, with  4,000 densely annotated. UniCalli employs asymmetric noising and a rasterized box map for spatial priors, trained on a mix of synthetic, labeled, and unlabeled data. The model achieves state-of-the-art generative quality with superior ligature continuity and layout fidelity, alongside stronger recognition. The framework successfully extends to other ancient scripts, including Oracle bone inscriptions and Egyptian hieroglyphs. Code and data can be viewed in this URL.
Figure 1: Generated by the UniCalli model, this image displays calligraphy from Li Bai’s poem ”Bring in the Wine”. Each column showcases a different master’s style to demonstrate the model’s versatility. Notably, especially in the Cursive script, the model generates contextually appropriate connecting strokes and character sizes based on adjacent characters. An English translation and the original Chinese text are provided in the lower corners. The complete works by each calligrapher are available in the Appendix G. (The calligraphic background has been manually edited for presentation.)
@@ 1 Introduction
Chinese calligraphy is a central vehicle of Chinese culture and a world heritage art form, practiced and studied by millions (Yangbo, 2022; Nihon Shuji, 2025). While recent deep learning has produced recognition (Luo et al., 2025; Zhou et al., 2025) and generative (Liao et al., 2023; 2024; Wu et al., 2020) systems, progress is hindered by scarce data and a long-tail distribution of styles. Consequently, existing generative methods are limited. On one hand, isolated-character synthesis and font transfer techniques (Zhang et al., 2024; Yang et al., 2024; Xie et al., 2021) can produce high-quality individual characters but ignore the holistic aesthetics of a finished work: page-level composition, spatial rhythm, and the crucial inter-character ligatures that convey artistic intent. On the other hand, general image generation models (Labs et al., 2025; Esser et al., 2024) and VLM-based systems (Bai et al., 2025; OpenAI, 2024; Gong et al., 2025) that attempt page-level synthesis often fail on correctness, rendering characters with improper forms or styles. This leaves a critical gap for generating complete calligraphic works that are both structurally sound and artistically coherent.
@@
  • A Large-Scale Annotated Calligraphy Dataset: We present a new, large-scale corpus of over 8,000 classical Chinese calligraphy works. More than 4,000 of these are annotated with script type, per-character bounding boxes, and modern transcriptions, providing a valuable resource to spur research in page-level analysis and generation.
  • UniCalli, a Unified Recognition and Generation Framework: We propose a novel diffusion transformer model that, for the first time, unifies page-level calligraphy generation and recognition. Its bidirectional attention mechanism enables globally coherent composition, moving beyond isolated characters to produce complete, stylistically consistent works.
@@ 2.1 Diffusion Models
Diffusion models (Ho et al., 2020; Song et al., 2020; Lipman et al., 2022) have recently emerged as a powerful class of generative models, achieving state-of-the-art results in high-fidelity image synthesis. Their core mechanism consists of two processes: a fixed forward process that gradually adds Gaussian noise to an input image until it becomes pure noise, and a learned reverse process. In the reverse process, a neural network, typically a U-Net (Rombach et al., 2022) or a Transformer (Peebles & Xie, 2023; Bao et al., 2023a), is trained to iteratively denoise a random input, step-by-step, to generate a clean sample.
@@ 2.2 Chinese Calligraphy Generation
Early studies in Chinese calligraphy generation, whether based on GANs (Wu et al., 2020; Xie et al., 2021; Tang & Lian, 2021) or high-fidelity diffusion models (Zhang et al., 2024; Liu & Lian, 2024; He et al., 2024; Dai et al., 2024), have primarily treated the task as one-shot or few-shot font generation. These methods excel at synthesizing isolated characters with accurate structure (Zeng et al., 2023) but do not address page-level composition.
@@ 3.3 Disentangling Style and Glyph via Conditional Dropout
During training, we observed that a standard conditional approach often leads the model to overfit to the styles of calligraphers from the long tail of the data distribution, thereby compromising correct glyph structure. To mitigate this issue, we propose a simple yet effective strategy to disentangle style from glyph formation through conditional dropout. During training, we stochastically replace the content condition with pure noise at a fixed probability $p_{\text{drop}}$ by setting its corresponding timestep $t_{c}$ to 1:
@@
Empirically, a dropout probability of $p_{\text{drop}}=0.05$ provides an effective balance. A higher probability tends to make the model disregard style information and generate more generic characters, while a lower probability can cause it to over-prioritize style, often resulting in abstract patterns with incoherent glyph structures.
@@ 4.1 Comparisons
Generation. We benchmark UniCalli’s generation capabilities against one-shot font generation method: FontDiffuser (Yang et al., 2024), VLM-based models: ChatGPT-5 (OpenAI, 2024), Ernie-4.5-Turbo (Research, 2025), and Doubao (Gong et al., 2025) in two settings. For reference-based synthesis, where a ground-truth image exists, visual comparisons are shown in Figure 3, and quantitative metrics (L1, SSIM, LPIPS, FID) are reported in Table 3. For reference-free synthesis, Figure 4 demonstrates UniCalli’s ability to generate a character in diverse styles given the same content. Finally, we conducted a user study (details in Appendix F) with 20 calligraphy enthusiasts who ranked the outputs based on Style Consistency, Glyph Accuracy, Naturalness, and Overall Preference, with results summarized in Table 1.
@@
Table 2: Character-level recognition accuracy on the held-out test set; bold indicates the best per row. “Doubao-1.5*” denotes Doubao-1.5-Thinking-Vision-Pro.
Category GPT-4o Ernie-4.5-Turbo-VL Qwen-2.5-VL-7B GOT-OCR2.0 $P$-OCRv5 Doubao-1.5* Ours
Cursive/Cao 0.073 0.255 0.127 0.091 0.091 0.200 0.109
Regular/Kai 0.502 0.616 0.600 0.314 0.396 0.588 0.688
Clerical/Li 0.293 0.453 0.453 0.187 0.160 0.507 0.518
Running/Xing 0.364 0.600 0.473 0.336 0.436 0.545 0.528
Seal/Zhuan 0.067 0.133 0.067 0.000 0.000 0.133 0.050
Total 0.380 0.534 0.482 0.266 0.324 0.510 0.540
@@
Table 4: Ablation study of our model’s key components. We start with a Baseline and incrementally add Joint Training, RoPE Duplication, and Conditional Dropout (Cond. Dropout).
@@ 4.2 Ablation Studies
Components ablation. We systematically evaluate the impact of the three main components of our model: (1) the Duplicate RoPE with Modulated Embedding, (2) the Joint Training Strategy utilizing synthetic, labeled, and unlabeled data, and (3) the Conditional Dropout mechanism. We start with a baseline model and incrementally add each component, measuring the performance at each stage. The quantitative results are summarized in Table 4.
@@
Table 5: Quantitative comparison of UniCalli and OracleNet on Oracle Bone Script tasks. Generation accuracy is evaluated by human experts and classified into three categories: Completely Correct, Largely Correct, and Completely Incorrect (detailed evaluation criteria are shown in Appendix E). In contrast, recognition accuracy is measured against deciphered ground truth. N/A denotes that a method is not applicable to the task.
@@
Oracle bone script. We trained our model on the HUST-OBC dataset (Wang et al., 2024) and structured our evaluation into three distinct tasks. First, we assessed the model’s ability to generate oracle bone script characters from corresponding modern Chinese characters. Second, we evaluated its performance on supervised oracle bone character recognition, using the Top-K accuracy metric to quantify precision. For the third task, we applied our model to a set of 100 undeciphered oracle bone characters to predict their potential modern Chinese character equivalents. The qualitative aspects of our study—namely, the generative task were evaluated by experts in oracle bone script from the School of History at Xiamen University, details in Appendix E, visual results are shown in Figure 6. Furthermore, we benchmarked our model against OracleNet (Zhou et al., 2025) on a curated set of 100 deciphered characters, as detailed in Table 5. Recognition accuracy was calculated directly, while generation correctness was assessed by experts.
Egyptian hieroglyphs. We conducted experiments on dataset (Umer, 2023) (see Appendix D for more details).
@@ C.3 Conditioning Signal Formulation
Appendix D Egyptian Hieroglyphs.
@@
Appendix E Human Expert Evaluation Criteria for Generation
@@
Appendix F User Study of General Calligraphy Generation
@@ F.1 Qualitative Feedback
Appendix G More Generation Results

### v2

@@ Abstract
Computational replication of Chinese calligraphy remains challenging. Existing methods falter, either creating high-quality isolated characters while ignoring column-level aesthetics like ligatures and spacing, or attempting column synthesis at the expense of calligraphic correctness. We introduce UniCalli, a unified diffusion framework for column-level recognition and generation. Training both tasks jointly is deliberate: recognition constrains the generator to preserve character structure, while generation provides style and layout priors. This synergy fosters concept-level abstractions that improve both tasks, especially in limited-data regimes. We curated a dataset of over 8,000 digitized pieces, with  4,000 densely annotated. UniCalli employs asymmetric noising and a rasterized box map for spatial priors, trained on a mix of synthetic, labeled, and unlabeled data. The model achieves state-of-the-art generative quality with superior ligature continuity and layout fidelity, alongside stronger recognition. The framework successfully extends to other ancient scripts: Oracle bone scripts and Egyptian hieroglyphs.
Figure 1: Generated by the UniCalli model, this image displays calligraphy from Li Bai’s poem ”Bring in the Wine”. Each column showcases a different master’s style to demonstrate the model’s versatility. Notably, especially in the Cursive script, the model generates contextually appropriate connecting strokes and character sizes based on adjacent characters. An English translation and the original Chinese text are provided in the lower corners. The complete works by each calligrapher are available in the Appendix I. (The calligraphic background has been manually edited for presentation.)
@@ 1 Introduction
Chinese calligraphy is a central vehicle of Chinese culture and a world heritage art form, practiced and studied by millions (Yangbo, 2022; Nihon Shuji, 2025). While recent deep learning has produced recognition (Luo et al., 2025; Zhou et al., 2025) and generative (Liao et al., 2023; 2024; Wu et al., 2020) systems, progress is hindered by scarce data and a long-tail distribution of styles. Consequently, existing generative methods are limited. On one hand, isolated-character synthesis and font transfer techniques (Zhang et al., 2024; Yang et al., 2024; Xie et al., 2021) can produce high-quality individual characters but ignore the holistic aesthetics of a finished work: column-level composition, spatial rhythm, and the crucial inter-character ligatures that convey artistic intent. On the other hand, general image generation models (Labs et al., 2025; Esser et al., 2024) and VLM-based systems (Bai et al., 2025; OpenAI, 2024; Gong et al., 2025) that attempt column-level synthesis often fail on correctness, rendering characters with improper forms or styles. This leaves a critical gap for generating complete calligraphic works that are both structurally sound and artistically coherent.
@@
  • A Large-Scale Annotated Calligraphy Dataset: We present a new, large-scale corpus of over 8,000 classical Chinese calligraphy works. More than 4,000 of these are annotated with script type, per-character bounding boxes, and modern transcriptions, providing a valuable resource to spur research in column-level analysis and generation.
  • UniCalli, a Unified Recognition and Generation Framework: We propose a novel diffusion transformer model that, for the first time, unifies column-level calligraphy generation and recognition. Its bidirectional attention mechanism enables globally coherent composition, moving beyond isolated characters to produce complete, stylistically consistent works.
@@ 2.1 Diffusion Models
Diffusion models (Ho et al., 2020; Song et al., 2020; Lipman et al., 2022) have recently emerged as a powerful class of generative models, achieving state-of-the-art results in high-fidelity image synthesis. Their core mechanism consists of two processes: a fixed forward process that gradually adds Gaussian noise to an input image until it becomes pure noise, and a learned reverse process. In the reverse process, a neural network, typically a U-Net (Rombach et al., 2022) or a Transformer (Peebles and Xie, 2023; Bao et al., 2023a), is trained to iteratively denoise a random input, step-by-step, to generate a clean sample.
@@ 2.2 Chinese Calligraphy Generation
Early studies in Chinese calligraphy generation, whether based on GANs (Wu et al., 2020; Xie et al., 2021; Tang and Lian, 2021) or high-fidelity diffusion models (Zhang et al., 2024; Liu and Lian, 2024; He et al., 2024; Dai et al., 2024), have primarily treated the task as one-shot or few-shot font generation. These methods excel at synthesizing isolated characters with accurate structure (Zeng et al., 2023) but do not address page-level composition.
@@ 3.3 Disentangling Style and Glyph via Conditional Dropout
Standard conditional training often leads to overfitting on long-tail stylistic data, where the model prioritizes rare style constraints at the expense of correct glyph structure. To mitigate this, we employ a strategy to disentangle style representation from glyph formation via conditional dropout. Specifically, we mask the content condition with pure noise at a probability $p_{\text{drop}}$ by setting the condition timestep $t_{c}$ to 1:
@@
Empirically, a dropout rate of $p_{\text{drop}}=0.05$ proves most effective. We observe that excessive dropout yields generic, style-agnostic outputs, while insufficient dropout leads to style over-prioritization, resulting in abstract patterns with incoherent glyph structures.
@@ 4.1 Comparisons
Generation. We benchmark UniCalli’s generation capabilities against one-shot font generation method: FontDiffuser (Yang et al., 2024), VLM-based models: ChatGPT-5 (OpenAI, 2024), Ernie-4.5-Turbo (Research, 2025), and Doubao (Gong et al., 2025) in two settings. For reference-based synthesis, where a ground-truth image exists, visual comparisons are shown in Figure 3, and quantitative metrics (L1, SSIM, LPIPS, FID) are reported in Table 3. For reference-free synthesis, Figure 4 demonstrates UniCalli’s ability to generate a character in diverse styles given the same content. Finally, we conducted a user study (details in Appendix H) with 20 calligraphy enthusiasts who ranked the outputs based on Style Consistency, Glyph Accuracy, Naturalness, and Overall Preference, with results summarized in Table 1.
@@
Table 2: Character-level recognition accuracy on the held-out test set; bold indicates the best per row. “Doubao-1.5*”, “Ernie-4.5*”, and “Qwen-2.5*” denote Doubao-1.5-Thinking-Vision-Pro, Ernie-4.5-Turbo-VL, and Qwen-2.5-VL-7B, respectively.
Category GPT-4o Ernie-4.5* Qwen-2.5* GOT-OCR2.0 $P$-OCRv5 Doubao-1.5* CalliReader Ours
Cursive/Cao 0.073 0.255 0.127 0.091 0.091 0.200 0.164 0.109
Regular/Kai 0.502 0.616 0.600 0.314 0.396 0.588 0.600 0.688
Clerical/Li 0.293 0.453 0.453 0.187 0.160 0.507 0.507 0.518
Running/Xing 0.364 0.600 0.473 0.336 0.436 0.545 0.658 0.528
Seal/Zhuan 0.067 0.133 0.067 0.000 0.000 0.133 0.067 0.050
Total 0.380 0.534 0.482 0.266 0.324 0.510 0.533 0.540
@@
Table 4: Ablation study of our model’s key components. Improved LPIPS and FID indicate enhanced diversity and realism, avoiding the mode collapse associated with pixel-wise averaging.
@@ 4.2 Ablation Studies
Components ablation. We systematically evaluate the impact of the three main components of our framework (Table 4): (1) the Duplicate RoPE with Modulated Embedding, (2) the Joint Training Strategy utilizing synthetic, labeled, and unlabeled data, and (3) the Conditional Dropout mechanism. To establish a rigorous comparison, we define the Baseline as a standard FLUX Labs et al. (2025) model trained exclusively on the generation task. In this baseline setup, the clean condition, noisy image, and mask are concatenated into a single input sequence processed by a single, continuous RoPE, without our proposed structural decoupling or joint optimization.
@@
Table 5: Quantitative comparison of UniCalli and OracleNet on Oracle Bone Script tasks. Generation accuracy is evaluated by human experts and classified into three categories: Completely Correct, Largely Correct, and Completely Incorrect (detailed evaluation criteria are shown in Appendix G). Recognition accuracy is measured against deciphered ground truth. N/A denotes that a method is not applicable to the task.
@@
Oracle bone script. We trained our model on the HUST-OBC dataset (Wang et al., 2024) and structured our evaluation into three distinct tasks. First, we assessed the model’s ability to generate oracle bone script characters from corresponding modern Chinese characters. Second, we evaluated its performance on supervised oracle bone character recognition, using the Top-K accuracy metric to quantify precision. For the third task, we applied our model to a set of 100 undeciphered oracle bone characters to predict their potential modern Chinese character equivalents. The qualitative aspects of our study—namely, the generative task were evaluated by experts in oracle bone script from the School of History at Xiamen University, details in Appendix G, visual results are shown in Figure 10. Furthermore, we benchmarked our model against OracleNet (Zhou et al., 2025) on a curated set of 100 deciphered characters, as detailed in Table 5. Recognition accuracy was calculated directly, while generation correctness was assessed by experts.
Egyptian hieroglyphs. We conducted experiments on dataset (Umer, 2023) (see Appendix F for more details).
@@ C.3 Conditioning Signal Formulation
C.4 Artifacts in Generated Samples
Origin of Visual Anomalies. A qualitative examination of certain generated samples (e.g., styles corresponding to Sun Guoting and Yan Zhenqing) reveals the presence of irregular ”white patches” or discontinuities within the stroke glyphs. We wish to clarify that these are not generative hallucinations or model failures, but rather high-fidelity reproductions of artifacts inherent to the training dataset.
Physical and Digital Causes. A significant portion of the training corpus is derived from historical stone stele rubbings (beike). These physical monuments have been subjected to centuries of natural weathering, erosion, and mechanical damage, resulting in surface pitting and cracks . When these historical rubbings undergo digital binarization for dataset curation, these physical imperfections—where the paper could not make contact with the eroded stone surface—are converted into binary ”noise.” Consequently, the model, which aims to approximate the underlying data distribution, learns to faithfully reconstruct these weathering artifacts as part of the stylistic texture.
Future Mitigation. We acknowledge that while historically authentic, these artifacts may detract from the aesthetic cohesion of the generated calligraphy for certain applications. In future work, we plan to implement rigorous data-level preprocessing, utilizing dedicated denoising algorithms to restore the integrity of the glyph strokes before training, thereby separating the calligraphic content from the historical degradation of the substrate.
Appendix D Evaluation on Public Benchmarks
To ensure a fair and robust evaluation of our recognizer, we extended our experiments to include third-party public datasets. However, it is important to note that, to the best of our knowledge, there are currently no publicly available datasets specifically dedicated to page-level Ancient Chinese Calligraphy. Existing public datasets in this domain primarily focus on isolated, individual characters Zhao et al. (2025).As the closest available alternative, we selected CalliBench Luo et al. (2025), a recently released benchmark for page-level calligraphy. It is crucial to emphasize that while CalliBench supports page-level evaluation, its data distribution differs significantly from our training objective: it predominantly consists of modern artistic fonts and contemporary calligraphy, rather than the historical ancient styles (e.g., Yan style, Cursive script) that UniCalli is optimized for. This introduces a substantial domain gap between the training and testing distributions.Despite this challenge, we evaluated UniCalli on the CalliBench ”Easy” subset against state-of-the-art general-purpose Vision-Language Models (VLMs) and specialized OCR systems. The results are reported in Table 6.
As shown in Table 6, UniCalli achieves a competitive F1 score of 0.498. While specialized OCR models like GOT-OCR2.0 achieve higher performance—likely due to their extensive training on broad, modern font datasets—UniCalli notably outperforms powerful general-purpose VLMs such as GPT-4o (F1: 0.400) and established industrial OCR systems like PP-OCRv5 (F1: 0.205). This result is encouraging: it demonstrates that although UniCalli was trained strictly on ancient historical data, it has learned robust, generalized features that allow for effective zero-shot transfer to unseen modern artistic styles.
Table 6: Quantitative comparison of recognition performance on the CalliBench Luo et al. (2025) (Easy subset). Note that UniCalli is trained exclusively on ancient historical calligraphy, whereas CalliBench consists largely of modern artistic fonts, introducing a domain gap.
Model Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ NED $\downarrow$
Ernie-4.5-Turbo-VL 0.542 0.481 0.482 0.637
Doubao-1.5-Thinking-Vision-Pro 0.442 0.513 0.462 0.655
$P$-OCRv5 0.372 0.291 0.205 0.859
GPT-4o 0.457 0.403 0.400 0.726
Qwen-2.5-VL-7B 0.440 0.736 0.534 0.710
GOT-OCR2.0 0.687 0.550 0.593 0.651
UniCalli 0.480 0.520 0.498 0.680
Appendix E Implementation Details of Unified Architecture
E.1 Zero-Shot Text Recognition via Feature Retrieval
Unlike traditional OCR systems that rely on autoregressive text decoding or classification heads, our unified framework performs recognition through a zero-shot, retrieval-based mechanism. This approach leverages the shared latent space between the generation and recognition tasks. The process consists of two phases:
  1. 1.
    Reference Library Construction (Offline):
    We construct a comprehensive reference feature library $\mathcal{L}$ prior to inference. We render images for all standard characters $c$ present in our vocabulary (derived from the standard .ttf font file used in training). Each rendered character image $I_{c}$ is passed through the pre-trained VAE encoder $\mathcal{E}$ to obtain its latent feature representation $\mathbf{z}_{c}$:
    $\mathcal{L}=\{(c,\mathbf{z}_{c})\mid\mathbf{z}_{c}=\mathcal{E}(I_{c}),\forall c\in\mathcal{V}\}$ (4)
    where $\mathcal{V}$ is the character vocabulary.
  2. 2.
    Inference and Matching:
    During the recognition task, the model outputs a predicted condition feature $\hat{\mathbf{z}}_{\text{pred}}$ (the tensor immediately preceding the VAE decoder). To determine the text content, we compute the cosine similarity between $\hat{\mathbf{z}}_{\text{pred}}$ and all reference features in $\mathcal{L}$. The recognized character $\hat{c}$ is determined by the nearest neighbor in this latent space:
    $\hat{c}=\operatorname*{argmax}_{c\in\mathcal{V}}\left(\frac{\hat{\mathbf{z}}_{\text{pred}}\cdot\mathbf{z}_{c}}{\|\hat{\mathbf{z}}_{\text{pred}}\|\|\mathbf{z}_{c}\|}\right)$ (5)
    This design avoids the need for an auxiliary OCR module and ensures that the recognition capability is intrinsically aligned with the generative features of the model.
E.2 Duplicate RoPE with Modulated Embedding
To process heterogeneous inputs (Image, Condition, and Mask) within a unified spatial context, we employ a Duplicate RoPE strategy augmented with Learnable Modulated Embeddings. This design ensures that the model can distinguish between input modalities without disrupting the spatial geometry defined by Rotary Positional Embeddings (RoPE).
Mechanism.
Let $\mathbf{X}\in\mathbb{R}^{L\times d}$ represent the input sequence of features for a specific modality $m\in\{\text{Image},\text{Condition},\text{Mask}\}$. We initialize a unique, learnable embedding vector $\mathbf{v}_{m}\in\mathbb{R}^{d}$ for each modality, initialized to zeros. This vector is broadcasted across the sequence length and added element-wise to the input features before the attention projection:
$\mathbf{X}^{\prime}_{m}=\mathbf{X}_{m}+\mathbf{v}_{m}$ (6)
Subsequently, the transformed features $\mathbf{X}^{\prime}_{m}$ are projected into Query ($\mathbf{Q}$) and Key ($\mathbf{K}$) states, to which RoPE is applied.
Preservation of Relative Positioning.
It is crucial to note that adding the modality embedding $\mathbf{v}_{m}$ does not interfere with the relative positional encoding properties of RoPE.
  • Spatial Structure: RoPE injects positional information via rotation based on the token index pospos.
  • Modality Identity: The learnable embedding $\mathbf{v}_{m}$ acts as a global, modality-specific “offset” or bias in the semantic space.
By superimposing $\mathbf{v}_{m}$, we allow the model to distinguish what the input is (e.g., “this is a condition token” vs. “this is an image token”) while RoPE simultaneously tells the model where the input is. Since the two signals operate orthogonally—one as a static semantic bias and the other as a dynamic frequency-based rotation—the underlying relative positional “grid” remains intact across all duplicated modalities.
Appendix F Egyptian Hieroglyphs.
@@ E.2 Duplicate RoPE with Modulated Embedding
Appendix G Human Expert Evaluation Criteria for Generation
@@
Appendix H User Study of General Calligraphy Generation
@@ H.1 Qualitative Feedback
Appendix I More Generation Results